OpenStudy (anonymous):

if a rectangle has on side as 12 and another as x. and a different rectangle has 2-x as one side and 16 as another, find the value of x so that the rectangles have the same area.

OpenStudy (anonymous):

12x = 16(2-x) 12x = 32 - 16x 28x = 32 x = 32/28

OpenStudy (anonymous):

\[\Delta _{1}=12(x)\] \[\Delta _{2}=16(2-x)\] If\[\Delta _{1}= \Delta _{2}\] 12(x)=16(2-x) 12x=32-16x 28x=32 x=32/28 =8/7